5(sqroot5)=25^(-10x-5)
Answer by stanbon(29602)

5(sqroot5)=25^(-10x-5)
5^(3/2) = 5^[2(-10x-5)]
3/2 = 2(-10x-5)
3 = 4(-10x-5)
3 = -40x-20
40x = -23
x = -23/40
===============
Cheers,
Stan H.